# Electron-phonon coupling induced topological phase transition in an $\alpha$-$T_{3}$ Haldane-Holstein model

### v1

@@ III.1 Bulk spectral properties
Let us consider the $\alpha=0.1$ case (Figs. 3$(a-c)$). As mentioned earlier, the mass term lifts the valley degeneracy. Also, the overall band spectrum is shifted vertically down as we increase $\lambda_{eph}$ further (the FB not being at $E=0$). Interestingly, at the two valleys, K and $\bf{K^{\prime}}$, the e-ph interaction makes the behaviour of the FBs contradictory (which otherwise looks symmetric when $\lambda_{eph}=0$ (see Fig. 2)), especially at $\lambda_{eph}=\lambda_{c}$ points. We clearly notice that in the $\lambda_{eph}<\lambda_{c}$ regime, the FB almost touches the CB at both the K and $\bf{K^{\prime}}$ valleys. Although, a prominent gap between VB and FB is maintained in the $\lambda_{eph}<\lambda_{c}$ regime. However, as soon as $\lambda_{eph}$ reaches a critical value, i.e. when $\lambda_{eph}=\lambda_{c}\approx 0.49$, the FB touches the VB at one of the valleys (K) and the band gap closes ($\Delta^{\textbf{K}}_{vf}=0$), while at the other valley ($\textbf{K}^{\prime}$), the spectrum remains gapped ($\Delta^{\textbf{K}}_{vf}\neq 0$). The band gap reopens and the gap persists if we increase $\lambda_{eph}$ further. Beyond $\lambda_{c}$ ($\lambda_{eph}>\lambda_{c}$), the behaviour of the FB is almost similar at both the valleys, especially for larger values of $\alpha$. Therefore, both in $\lambda_{eph}<\lambda_{c}$ and $\lambda_{eph}>\lambda_{c}$ regimes, the spectrum remains gapped, implying it to be an insulator, and at $\lambda_{eph}=\lambda_{c}$ the bands touch, signifying a semi-metallic ($SM$) behaviour.
@@
We need to compute the topological properties for different $\lambda_{eph}$ regimes to confirm the topological nature of the phase, which we shall show in the later sections (III.2 and 15). This phenomenon of band closing and opening at the Dirac points may give rise to a topological phase transition that is solely caused by tuning the e-ph interaction strength. This is the central result of the paper. Smaller values of $\alpha$ in the range [0.1:0.3] demonstrate similar behaviour (Fig. 3) with different $\lambda_{c}$’s (listed in Table 2). The intermediate $\alpha$-cases ($0.4\leq\alpha\leq 0.6$) are shown in Fig. 4 where we observe the same phenomena, except that one notices for $\alpha=0.6$ case, the FB and the VB nearly touch each other even when $\lambda_{eph}<\lambda_{c}$ (the values of $\lambda_{c}$’s are mentioned in Table 2) region. This feature persists for larger values of $\alpha$ ($\alpha>0.6$) and it needs to be addressed carefully. To do so, we plot the band structure in Fig. 5 for $\alpha=0.7$, where it is clearly shown that the VB and the FB touch each other below a certain critical value, namely, $\lambda_{c_{1}}\approx 0.28$ which may describe a $SM$ phase in the $\lambda_{eph}<\lambda_{c_{1}}$ regime, and will hold even when $\lambda_{eph}=0$. Then, in the vicinity of $\lambda_{eph}\approx\lambda_{c_{1}}$, the gap between VB and FB opens for the first time (can be seen clearly if we zoom in), signalling an insulating behaviour, and the gap stays intact up to a second critical point, namely, $\lambda_{c_{2}}\approx 0.43$, at which the gap closes, referring a re-onset of a semi-metallic phase. The bulk spectrum is gapped beyond $\lambda_{c_{2}}$. For $\alpha=0.8$ and $\alpha=0.9$, the scenario is a bit more interesting. Unlike $\alpha=0.7$, for $\alpha=0.8$, we observe in Fig. 6$(a-b)$ that the bulk bands remain gapped (can be seen clearly if we zoom in) in $\lambda_{eph}<\lambda_{c_{1}}$ regime (including $\lambda_{eph}=0$), signifying an insulating (not $SM$ as for $\alpha=0.7$) phase till $\lambda_{eph}=\lambda_{c_{1}}\approx 0.2$, where the FB and VB touch each other for the first time (Fig. 6$(c)$) and the insulating to $SM$ transition takes place. As we tune $\lambda_{eph}$ above $\lambda_{c_{1}}$, we observe the same phenomena as it is shown for $\alpha=0.7$ case (Fig. 5), that is, in $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ regime (Fig. 6$(d)$), the FB and VB remain gapped (can be seen clearly if we zoom in) denoting an insulating phase till $\lambda_{eph}=\lambda_{c_{2}}\approx 0.44$ at which the system again shows a $SM$ (Fig. 6$(e)$) nature and for $\lambda_{eph}>\lambda_{c_{2}}$ (Fig. 6$(f)$) it behaves like an insulator, alike it does for $\alpha=0.7$ case. Similar observations hold for $\alpha=0.9$ (not shown here). We wish to mention that for $10.8\leq\alpha<$ 1, this feature of multiple phase transition becomes more prominent as we approach the dice lattice ($\alpha=1$). Therefore, for higher $\alpha$-cases ($\alpha>0.6$), we encounter two situations, one is for $0.80.6<\alpha<$ 0.8 and another is for $10.8\leq\alpha<$ 1. In the former case, we get two critical $\lambda_{c}$-points, namely $\lambda_{c_{1}}$ and $\lambda_{c_{2}}$, below (even when $\lambda_{eph}=0$) and above which the system respectively remains semi-metallic and insulating respectively. In between $\lambda_{c_{1}}$ and $\lambda_{c_{2}}$, it behaves like an insulator. So, the system undergoes an $SM$-insulator-$SM$-insulator transition in the former case, while in the latter case, the system inherits an insulator-$SM$-insulator-$SM$-insulator transition. The nature of the gap (topological or trivial), will be ascertained in Sec. III.2 and Sec. 15. Hence, the band topology in our study is substantially modified by the polaron formation, which is governed by two factors: the renormalized amplitudes $\tilde{t}$ and $\tilde{\lambda}$ (Eq. (8)) and also the interplay between $MS_{z}$ and $\lambda_{eph}^{2}\hbar\omega_{0}$ (last two terms of Eq. (9)). The former causes the band narrowing and the latter is responsible for the competitive effects between the mass term and the polaron shift energy. Moreover, these polaronic markers make the variations of the band spectra (especially those of the FB and the VBs) different for different ranges of $\alpha$. Specifically, for higher values of $\alpha$, the correlation between $M$, $\alpha$ and $\lambda_{eph}$ becomes stronger, giving rise to multiple phase transitions. In the case of the dice lattice ($\alpha=1$), the flat band remains flat without any distortion and there is no occurrence of band gap closing phenomena for any values of the e-ph coupling $\lambda_{eph}$ (not shown here). So, no $\lambda_{eph}$ yields a topological phase transition. It is worth mentioning that the values of $\lambda_{c}$ are different for different $\alpha$ cases (see Table 2 and Table 2), which ensures that we shall have a phase transition for all $\alpha$ values between $\alpha=0$ to $1$, albeit with different $\lambda_{c}$ values.
@@
III.2 Edge modes of a semi-infinite ${\boldsymbol{\alpha\--T_{3}}}$ ribbon
@@
Figure 7: Energy spectra (in units of $t$) of the edge states are shown for a zigzag edged semi-infinite ribbon of $\alpha=0.1$ for $(a)$ $\lambda_{eph}=0.3$, $(b)$ $\lambda_{eph}=0.49$, and $(c)$ $\lambda_{eph}=0.6$, and of $\alpha=0.2$ for $(d)$ $\lambda_{eph}=0.3$, $(e)$ $\lambda_{eph}=0.48$, and $(f)$ $\lambda_{eph}=0.6$. Other parameters are the same as those in Fig. 3.
@@
We begin by referring to Fig. 7, where we show the edge states for lower $\alpha$-values ($\alpha=0.1$ and $\alpha=0.2$ marked on the right edge). As stated above in Sec. III.1, the bulk gap closes at a critical $\lambda_{c}$ and it remains gapped corresponding to $\lambda_{eph}<\lambda_{c}$ and $\lambda_{eph}>\lambda_{c}$. We wish to ascertain the existence of edge states that distinguishes a topologically nontrivial phase from a trivial one in both scenarios. Below a critical $\lambda_{c}$, Figs. 7 $(a)$ and $(d)$ display a prominent set of edge states traversing from CB to VB through FB (and vice versa) for $\alpha=0.1$ and $\alpha=0.2$ respectively for $\lambda_{eph}<\lambda_{c}$. We notice that a pair of edge states emerge from different valleys in the bulk, gather at the FBs and hence cross over to the CBs. By looking at the slope of the edge states, that is, $\partial E/\partial k$, which is a measure of the velocity of the electron, we infer that the flow of the edge currents is counterpropagating, as it should be. These edge states are the chiral edge states of a Chern insulator, appearing in the regime of $\lambda_{eph}<\lambda_{c}$. The nature of the edge states for the $\alpha=0.2$ case are distinct, in the sense that they are crossing the FB at different points. It is also visible in Figs. 7 $(b)$ and $(e)$ that these chiral edges persist up to $\lambda_{eph}\approx\lambda_{c}$ and disappear beyond that. These are presented in Figs. 7 $(c)$ and $(f)$, that for values above $\lambda_{c}$, the edge states completely disappear and bulk spectra become gapped, signifying the transition of the system to a trivial phase. The critical values of $\lambda_{eph}$ corresponding to the transitions for the $\alpha=0.1$ and $\alpha=0.2$ cases are listed in Table 2. Therefore, in $\alpha$-$T_{3}$ systems (with smaller $\alpha$ values), one can generate topological insulating phases via only tuning $\lambda_{eph}$ for a particular value of $\alpha$ below a certain $\lambda_{c}$, beyond which the system goes into a trivial insulating phase.
@@
Figure 8: Energy spectra (in units of $t$) of the edge states are shown for a zigzag edged semi-infinite ribbon of $\alpha=0.7$ for $(a)$ $\lambda_{eph}=0.2$, $(b)$ $\lambda_{eph}=\lambda_{c_{1}}=0.28$, $(c)$ $\lambda_{eph}=0.35$, $(d)$ $\lambda_{eph}=\lambda_{c_{2}}=0.43$, and $(e)$ $\lambda_{eph}=0.6$. Other parameters are the same as those in Fig. 3.
@@
Figure 9: Energy spectra (in units of $t$) of the edge states are shown for a zigzag edged semi-infinite ribbon of $\alpha=0.8$ for $(a)$ $\lambda_{eph}=0.0$, $(b)$ $\lambda_{eph}=0.15$, $(c)$ $\lambda_{eph}=\lambda_{c_{1}}=0.28$, $(d)$ $\lambda_{eph}=0.35$, $(e)$ $\lambda_{eph}=\lambda_{c_{2}}=0.44$, and $(f)$ $\lambda_{eph}=0.6$. Other parameters are the same as those in Fig. 3.
@@
Next, let us study the characteristics of the edge states for higher $\alpha$ values, and as a specific case, consider $\alpha=0.7$, presented in Fig. 8. In reference to its bulk properties displayed in Fig. 5, we shall examine the edge states for different regimes of $\lambda_{eph}$. As discussed in Fig. 5$(a)$, the bulk FB and VB remain in contact with each other for the $\lambda_{eph}\lesssim\lambda_{c_{1}}$ region. We notice its signature in Fig. 8$(a)$, where a pair of counterpropagating edge states emerge near each $\bf{K}$-valley, passing through the FB for the $\lambda_{eph}\lesssim\lambda_{c_{1}}$ regime. However, in this regime of $\lambda_{eph}$, the notion of edge states is not important as the system does not have any bulk gap, inferring it to be a usual semi-metal. The edge states connecting the VB and CB through the FB are gapped till $\lambda_{eph}\approx\lambda_{c_{1}}$ at which the edge states at one $\bf{K}$-valley touch for the first time (see Fig. 8$(b)$), thereby generating a conducting edge mode. In the intermediate region (see Fig. 8$(c)$), i.e. for $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$, the system clearly exhibits the presence of edge states indicating a topologically nontrivial (Chern insulating) phase. It is evident in Fig.  8$(c)$ that the edge states are counterpropagating, and they cross the FB at the two edges for $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$. Therefore, for $\alpha=0.7$, there seems to be a re-entrant mechanism to the $SM$ phase, which may be achieved entirely by tuning the e-ph coupling strength. Around $\lambda_{eph}\approx\lambda_{c_{2}}$, the edge states start fading out (shown in Fig. 8$(d)$) at one $\bf{K}$-valley and is completely disappear above $\lambda_{c_{2}}$ (see Fig. 8 $(e)$). Undoubtedly, the $\lambda_{eph}>\lambda_{c_{2}}$ region refers to a trivial insulator with no sign of edge states. As suggested in the discussion of the bulk spectra (Sec. III.1) that multiple phase transitions (insulator-$SM$-insulator-$SM$-insulator) can occur for $\alpha>0.7$, we explicitly plot the edge states for $\alpha=0.8$ in Fig. 9 which ascertains whether the insulating phases are topological. In Fig. 9$(a)$ we notice that in the absence of e-ph coupling, a pair of prominent edge states cross the FB at $\bf{K}$ or $\bf{K^{\prime}}$ valley, signifying a topologically nontrivial Chern insulating phase (unlike for $\alpha=0.7$ where it is $SM$), which remain intact in the $\lambda_{eph}<\lambda_{c_{1}}$ regime (Fig.  9$(b)$) till $\lambda_{eph}=\lambda_{c_{1}}$. Beyond this value one pair of edge states becomes gapped at one valley, while in the other valley it remains gapless (Fig. 9$(c)$). However, such as for the $\alpha>0.7$ case, it is vividly seen in Figs. 9$(d-f)$ that the counterpropagating edge states resurface in the $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ regime, persist up to $\lambda_{eph}=\lambda_{c_{2}}$ and completely vanish beyond $\lambda_{c_{2}}$. So, for $\alpha=0.8$ as well, the re-entrant scenario to the $SM$ phase still holds (also true for $\alpha=0.9$, not shown here). To confirm that the edge modes indeed correspond to a Chern insulating phase, we compute the topological properties and discuss them for each of the regions of $\lambda_{eph}$ (as indicated above) in the following section (Sec. 15). The phase transition points, namely $\lambda_{c_{1}}$ and $\lambda_{c_{2}}$ for $\alpha=0.7,0.8$ and $0.9$ are listed in Table 2.

### v2

@@ III.1 Bulk spectral properties
Let us consider the $\alpha=0.1$ case (Figs. 3$(a-c)$). As mentioned earlier, the mass term lifts the valley degeneracy. Also, the overall band spectrum is shifted vertically down as we increase $\lambda_{eph}$ further (the FB not being at $E=0$). Interestingly, at the two valleys, K and $\bf{K^{\prime}}$, the e-ph interaction makes the behaviour of the FBs contradictory (which otherwise looks symmetric when $\lambda_{eph}=0$ (see Fig. 2)), especially at $\lambda_{eph}=\lambda_{c}$ points. We clearly notice that in the $\lambda_{eph}<\lambda_{c}$ regime (Fig. 3$(a)$), the FB almost touches the CB at both the K and $\bf{K^{\prime}}$ valleys. Although, a prominent gap between VB and FB is maintained in the $\lambda_{eph}<\lambda_{c}$ regime. However, as soon as $\lambda_{eph}$ reaches a critical value, i.e. when $\lambda_{eph}=\lambda_{c}=0.49$, the FB touches the VB (Fig. 3$(b)$) at one of the valleys (K) and the band gap closes ($\Delta^{\textbf{K}}_{vf}=0$), while at the other valley ($\textbf{K}^{\prime}$), the spectrum remains gapped ($\Delta^{\textbf{K}}_{vf}\neq 0$). The band gap re-opens and the gap persists if we increase $\lambda_{eph}$ further. Beyond $\lambda_{c}$ ($\lambda_{eph}>\lambda_{c}$), the behaviour of the FB is almost similar at both the valleys (Fig. 3$(c)$), especially for larger values of $\alpha$. Therefore, both in $\lambda_{eph}<\lambda_{c}$ and $\lambda_{eph}>\lambda_{c}$ regimes, the spectrum remains gapped, implying it to be an insulator, and at $\lambda_{eph}=\lambda_{c}$ the bands touch, signifying a semi-metallic ($SM$) behaviour. We need to compute the topological properties for different $\lambda_{eph}$ regimes to confirm the topological nature of the phase, which we shall show in the later sections (III.2 and 15). This phenomenon of band closing and opening at the Dirac points may give rise to a topological phase transition that is solely caused by tuning the e-ph interaction strength. This is the central result of the paper. Smaller values of $\alpha$ in the range $[0.1:0.3]$ demonstrate similar behaviour (Fig. 3) with different $\lambda_{c}$’s (listed in Table 2).
@@
The intermediate $\alpha$-cases ($0.4\leq\alpha\leq 0.6$) are shown in Fig. 4 where we observe the same phenomena, except that one notices for $\alpha=0.6$ case, the FB and the VB nearly touch each other even when $\lambda_{eph}<\lambda_{c}$ (the values of $\lambda_{c}$ are mentioned in Table 2) region (can be seen clearly if we zoom in Fig. 4$(g)$). This feature persists for larger values of $\alpha$ ($\alpha>0.6$) and it needs to be addressed carefully. To do so, we plot the band structure in Fig. 5 for $\alpha=0.7$, where it is clearly shown that the VB and the FB touch each other below a certain critical value, namely, $\lambda_{c_{1}}=0.28$ which may describe a $SM$ phase in the $\lambda_{eph}<\lambda_{c_{1}}$ regime (Fig. 5$(b)$), and will hold even when $\lambda_{eph}=0$ (Fig. 5$(a)$). Then, in the vicinity of $\lambda_{eph}=\lambda_{c_{1}}$, the gap between VB and FB opens for the first time (can be seen clearly if we zoom in Fig. 5$(c)$), signalling an insulating behaviour, and the gap stays intact in the $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ regime (can be seen clearly if we zoom in Fig. 5$(d)$) up to a second critical point, namely, $\lambda_{c_{2}}=0.43$, at which the gap closes (Fig. 5$(e)$), referring a re-onset of a $SM$ phase. The bulk spectrum is gapped beyond $\lambda_{c_{2}}$ (Fig. 5$(f)$). For $\alpha=0.8$ and $\alpha=0.9$, the scenario is a bit more interesting. Unlike $\alpha=0.7$, for $\alpha=0.8$, we observe in Fig. 6$(a-b)$ that the bulk bands remain gapped (can be seen clearly if we zoom in) in $\lambda_{eph}<\lambda_{c_{1}}$ regime (including $\lambda_{eph}=0$), signifying an insulating (not $SM$ as for $\alpha=0.7$) phase till $\lambda_{eph}=\lambda_{c_{1}}=0.2$, where the FB and VB touch each other for the first time (Fig. 6$(c)$) and the insulating to $SM$ transition takes place. As we tune $\lambda_{eph}$ above $\lambda_{c_{1}}$, we observe the same phenomena as it is shown for $\alpha=0.7$ case (Fig. 5), that is, in $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ regime (Fig. 6$(d)$), the FB and VB remain gapped (can be seen clearly if we zoom in) denoting an insulating phase till $\lambda_{eph}=\lambda_{c_{2}}=0.44$ at which the system again shows a $SM$ (Fig. 6$(e)$) nature and for $\lambda_{eph}>\lambda_{c_{2}}$ (Fig. 6$(f)$) it behaves like an insulator, alike it does for $\alpha=0.7$ case. Similar observations hold for $\alpha=0.9$ (not shown here). The values of $\lambda_{c_{1}}$ and $\lambda_{c_{2}}$ for $0.7$, $0.8$, and $0.9$ are listed in Table 2. We wish to mention that for $10.8\leq\alpha<$ 1, this feature of multiple phase transition becomes more prominent as we approach the dice lattice ($\alpha=1$). Therefore, for higher $\alpha$-cases ($\alpha>0.6$), we encounter two situations, one is for $0.80.6<\alpha<$ 0.8 and another is for $10.8\leq\alpha<$ 1. In the former case, we get two critical $\lambda_{c}$-points, namely $\lambda_{c_{1}}$ and $\lambda_{c_{2}}$, below (even when $\lambda_{eph}=0$) and above which the system respectively remains semi-metallic and insulating respectively. In between $\lambda_{c_{1}}$ and $\lambda_{c_{2}}$, it behaves like an insulator. So, the system undergoes an $SM$-insulator-$SM$-insulator transition in the former case, while in the latter case, the system inherits an insulator-$SM$-insulator-$SM$-insulator transition. The nature of the gap (topological or trivial), will be ascertained in Sec. III.2 and Sec. 15. Hence, the band topology in our study is substantially modified by the polaron formation, which is governed by two factors: the renormalized amplitudes $\tilde{t}$ and $\tilde{\lambda}$ (Eq. (8)) and also the interplay between $MS_{z}$ and $\lambda_{eph}^{2}\hbar\omega_{0}$ (last two terms of Eq. (9)). The former causes the band narrowing and the latter is responsible for the competitive effects between the mass term and the polaron shift energy. Moreover, these polaronic markers make the variations of the band spectra (especially those of the FB and the VBs) different for different ranges of $\alpha$. Specifically, for higher values of $\alpha$, the correlation between $M$, $\alpha$ and $\lambda_{eph}$ becomes stronger, giving rise to multiple phase transitions. In the case of the dice lattice ($\alpha=1$), the flat band remains flat without any distortion and there is no occurrence of band gap closing phenomena for any values of the e-ph coupling $\lambda_{eph}$ (not shown here). So, no $\lambda_{eph}$ yields a topological phase transition. It is worth mentioning that the values of $\lambda_{c}$ are different for different $\alpha$ cases (see Table 2 and Table 2), which ensures that we shall have a phase transition for all $\alpha$ values between $\alpha=0$ to $1$, albeit with different $\lambda_{c}$ values.
@@
III.2 Edge modes of a semi-infinite $\alpha$-$T_{3}$ ribbon
@@
Figure 7: Energy spectra (in units of $t$) of the edge states are shown for a zigzag edged semi-infinite ribbon as a function of dimensionless momenta, $k_{x}$ (multiplied by the lattice constant) of $\alpha=0.1$ for $(a)$ $\lambda_{eph}=0.3$ ($\lambda_{eph}<\lambda_{c}$), $(b)$ $\lambda_{eph}=\lambda_{c}=0.49$, and $(c)$ $\lambda_{eph}=0.6$ ($\lambda_{eph}>\lambda_{c}$), and of $\alpha=0.2$ for $(d)$ $\lambda_{eph}=0.3$ ($\lambda_{eph}<\lambda_{c}$), $(e)$ $\lambda_{eph}=\lambda_{c}=0.48$, and $(f)$ $\lambda_{eph}=0.6$ ($\lambda_{eph}>\lambda_{c}$). Other parameters are the same as those in Fig. 3. The values of $\lambda_{c}$ are mentioned in Table 2.
@@
We begin by referring to Fig. 7, where we show the edge states for lower $\alpha$-values ($\alpha=0.1$ and $\alpha=0.2$ marked on the right edge). As stated above in Sec. III.1, the bulk gap closes at a critical $\lambda_{c}$ and it remains gapped corresponding to $\lambda_{eph}<\lambda_{c}$ and $\lambda_{eph}>\lambda_{c}$. We wish to ascertain the existence of edge states that distinguishes a topologically nontrivial phase from a trivial one in both scenarios. Below a critical $\lambda_{c}$, Figs. 7 $(a)$ and $(d)$ display a prominent set of edge states traversing from CB to VB through FB (and vice versa) for $\alpha=0.1$ and $\alpha=0.2$, respectively for the $\lambda_{eph}<\lambda_{c}$ regime. We notice that a pair of edge states emerge from different valleys in the bulk, gather at the FBs and hence cross over to the CBs. By looking at the slope of the edge states, that is, $\partial E/\partial k$, which is a measure of the velocity of the electron, we infer that the flow of the edge currents is counterpropagating, as it should be. These edge states are the chiral edge states of a Chern insulator, appearing in the regime of $\lambda_{eph}<\lambda_{c}$. The nature of the edge states for the $\alpha=0.2$ case are distinct, in the sense that they are crossing the FB at different points. It is also visible in Figs. 7 $(b)$ and $(e)$ that these chiral edges persist up to $\lambda_{eph}=\lambda_{c}$ and disappear beyond that. These are presented in Figs. 7 $(c)$ and $(f)$, that for values above $\lambda_{c}$, the edge states completely disappear and bulk spectra become gapped, signifying the transition of the system to a trivial phase. The critical values of $\lambda_{eph}$ corresponding to the transitions for the $\alpha=0.1$ and $\alpha=0.2$ cases are listed in Table 2. Therefore, in $\alpha$-$T_{3}$ systems (with smaller $\alpha$ values), one can generate topological insulating phases via only tuning $\lambda_{eph}$ for a particular value of $\alpha$ below a certain $\lambda_{c}$, beyond which the system goes into a trivial insulating phase.
@@
Figure 8: Energy spectra (in units of $t$) of the edge states are shown for a zigzag edged semi-infinite ribbon as a function of dimensionless momenta, $k_{x}$ (multiplied by the lattice constant) of $\alpha=0.7$ for $(a)$ $\lambda_{eph}=0$, $(b)$ $\lambda_{eph}=0.2$ ($\lambda_{eph}<\lambda_{c_{1}}$), $(c)$ $\lambda_{eph}=\lambda_{c_{1}}=0.28$, $(d)$ $\lambda_{eph}=0.35$ ($\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$), $(e)$ $\lambda_{eph}=\lambda_{c_{2}}=0.43$, and $(f)$ $\lambda_{eph}=0.6$ ($\lambda_{eph}>\lambda_{c_{2}}$). Other parameters are the same as those in Fig. 5. The values of $\lambda_{c_{1}}$ and $\lambda_{c_{2}}$ are mentioned in Table 2.
@@
Figure 9: Energy spectra (in units of $t$) of the edge states are shown for a zigzag edged semi-infinite ribbon as a function of dimensionless momenta, $k_{x}$ (multiplied by the lattice constant) of $\alpha=0.8$ for $(a)$ $\lambda_{eph}=0$, $(b)$ $\lambda_{eph}=0.15$ ($\lambda_{eph}<\lambda_{c_{1}}$), $(c)$ $\lambda_{eph}=\lambda_{c_{1}}=0.2$, $(d)$ $\lambda_{eph}=0.35$ ($\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$), $(e)$ $\lambda_{eph}=\lambda_{c_{2}}=0.44$, and $(f)$ $\lambda_{eph}=0.6$ ($\lambda_{eph}>\lambda_{c_{2}}$). Other parameters are the same as those in Fig. 6. The values of $\lambda_{c_{1}}$ and $\lambda_{c_{2}}$ are mentioned in Table 2.
@@
Next, let us study the characteristics of the edge states for higher $\alpha$ values, and as a specific case, consider $\alpha=0.7$, presented in Fig. 8. In reference to its bulk properties displayed in Fig. 5, we shall examine the edge states for different regimes of $\lambda_{eph}$. As discussed in Figs. 5$(a-b)$, the bulk FB and VB remain in contact with each other for the $\lambda_{eph}\lesssim\lambda_{c_{1}}$ region, we notice its signature in Figs. 8$(a-b)$, where a pair of counterpropagating edge states emerge near each $\bf{K}$-valley, passing through the FB for the $\lambda_{eph}\lesssim\lambda_{c_{1}}$ regime. However, in this regime of $\lambda_{eph}$, the notion of edge states is not important as the system does not have any bulk gap, inferring it to be a usual semi-metal. The edge states connecting the VB and CB through the FB are gapped till $\lambda_{eph}=\lambda_{c_{1}}$ at which the edge states at one $\bf{K}$-valley touch for the first time (can be seen clearly if we zoom in Fig. 8$(c)$), thereby generating a conducting edge mode. In the intermediate region (see Fig. 8$(d)$), i.e. for $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$, the system clearly exhibits the presence of edge states indicating a topologically nontrivial (Chern insulating) phase. It is evident in Fig. 8$(d)$ that the edge states are counterpropagating, and they cross the FB at the two edges for $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$. Therefore, for $\alpha=0.7$, there seems to be a re-entrant mechanism to the $SM$ phase, which may be achieved entirely by tuning the e-ph coupling strength. Around $\lambda_{eph}=\lambda_{c_{2}}$, the edge states start fading out (shown in Fig. 8$(e)$) at one $\bf{K}$-valley and is completely disappear above $\lambda_{c_{2}}$ (see Fig. 8$(f)$). Undoubtedly, the $\lambda_{eph}>\lambda_{c_{2}}$ region refers to a trivial insulator with no sign of edge states. As suggested in the discussion of the bulk spectra (Sec. III.1) that multiple phase transitions (insulator-$SM$-insulator-$SM$-insulator) can occur for $\alpha>0.7$, we explicitly plot the edge states for $\alpha=0.8$ in Fig. 9 which ascertains whether the insulating phases are topological. In Fig. 9$(a)$ we notice that in the absence of e-ph coupling, a pair of prominent edge states cross the FB at $\bf{K}$ or $\bf{K^{\prime}}$ valley, signifying a topologically nontrivial Chern insulating phase (unlike for $\alpha=0.7$ where it is $SM$), which remain intact in the $\lambda_{eph}<\lambda_{c_{1}}$ regime (Fig.  9$(b)$) till $\lambda_{eph}=\lambda_{c_{1}}$. At this value one pair of edge states becomes gapped at one valley, while in the other valley it remains gapless (Fig. 9$(c)$). However, such as for the $\alpha>0.7$ case, it is vividly seen in Figs. 9$(d-f)$ that the counterpropagating edge states resurface in the $\lambda_{c_{1}}<\lambda_{eph}<\lambda_{c_{2}}$ regime, persist up to $\lambda_{eph}=\lambda_{c_{2}}$ and completely vanish beyond $\lambda_{c_{2}}$. So, for $\alpha=0.8$ as well, the re-entrant scenario to the $SM$ phase still holds (also true for $\alpha=0.9$, not shown here). To confirm that the edge modes indeed correspond to a Chern insulating phase, we compute the topological properties and discuss them for each of the regions of $\lambda_{eph}$ (as indicated above) in the following section (Sec. 15). The phase transition points, namely $\lambda_{c_{1}}$ and $\lambda_{c_{2}}$ for $\alpha=0.7,0.8$ and $0.9$ are listed in Table 2.